# VIRTUE: Visual-Interactive Text-Image Universal Embedder

### v1

@@ E.3 Reevaluating MMEB with Further Finetuning on SCaR-train
To evaluate the impact of incorporating SCaR-train on MMEB performance, we reevaluate VLM2Vec-2B, VLM2Vec-7B, MMRet-7B, UniME-7B, VIRTUE-2B, and VIRTUE-7B with SCaR-train (+SCaR-train) (Sec. 5.3) and compare them against the original checkpoints trained solely on MMEB-train. As shown in Tab. 7, all baselines exhibit a degraded performance, likely because the models overfit or shift focus toward the SCaR distribution for incremental learning (Li & Hoiem, 2016). Notably, VLM2Vec and MMRet suffer severe degradation, suggesting stronger susceptibility to catastrophic forgetting (French, 1999) when their training relies more heavily on additional data. In contrast, UniME, with its multi-stage contrastive learning design, demonstrates greater robustness against forgetting. Meanwhile, VIRTUE maintains a performance comparable to the original MMEB-trained checkpoints, with VIRTUE-2B even achieving a slight 0.1-point gain. We attribute this stability to the segmentation streamline, which enables the use of optional visual prompts and makes VIRTUE more adaptable as a universal embedder capable of handling both visual-interactive and non-visual-interactive tasks.
@@ E.5 Qualitative Results on SCaR
E.6 Detailed Scores of MMEB
@@
![[Uncaptioned image]](Figures/scar-examples/refcoco_plus.png)
![[Uncaptioned image]](Figures/scar-examples/refcocog.png)
![[Uncaptioned image]](Figures/scar-examples/visualgenome.png)
![[Uncaptioned image]](Figures/scar-examples/coco-stuff.png)
![[Uncaptioned image]](Figures/scar-examples/ADE20K.jpg)

### v2

@@ E.3 Reevaluating MMEB with Further Finetuning on SCaR-train
To evaluate the impact of incorporating SCaR-train on MMEB performance, we reevaluate VLM2Vec-2B, VLM2Vec-7B, MMRet-7B, UniME-7B, VIRTUE-2B, and VIRTUE-7B with SCaR-train (+SCaR-train) (Sec. 5.3) and compare them against the original checkpoints trained solely on MMEB-train. As shown in Tab. 10, all baselines exhibit a degraded performance, likely because the models overfit or shift focus toward the SCaR distribution for incremental learning (Li and Hoiem, 2016). Notably, VLM2Vec and MMRet suffer severe degradation, suggesting stronger susceptibility to catastrophic forgetting (French, 1999) when their training rely more heavily on additional data. In contrast, UniME, with its multi-stage contrastive learning design, demonstrates greater robustness against forgetting. Meanwhile, VIRTUE maintains a performance comparable to the original MMEB-trained checkpoints, with VIRTUE-2B even achieving a slight 0.1-point gain. We attribute this stability to the segmentation streamline, which enables the use of optional visual prompts and makes VIRTUE more adaptable as a universal embedder capable of handling both visual-interactive and non-visual-interactive tasks.
@@ E.5 Qualitative Results on SCaR
E.6 Analysis on Latency and Memory
| Query Mode | Time (per sample) | Memory | Performance | |
| I2T (SCaR) | ||||
| VIRTUE-7B | w/ visual prompts | 422ms | 18.8GB | 27.8 |
| w/o visual prompts (uniform points) | 427ms | 18.8GB | 16.1 | |
| Cropping | 139ms | 17.7GB | 22.8 | |
| Text-only | 340ms | 17.7GB | 23.9 | |
| I2I (Retrieval and Visual Grounding on MMEB) | ||||
| VIRTUE-7B | w/ uniform points | 402ms | 18.8GB | 70.7 |
| w/o uniform points | 357ms | 17.7GB | 46.5 | |
| VLM2Vec-7B | N/A | 406ms | 18.6GB | 67.9 |
| MMRet-7B | N/A | 538ms | 29.0GB | 75.8 |
| UniME-7B | N/A | 419ms | 28.9GB | 66.4 |
We report the average end-to-end inference time (ms) and GPU memory (GB) for four distinct operational modes of VIRTUE-7B on the SCaR benchmark (I2T) and provide baselines (VLM2Vec, MMRet, and UniME) for I2I (retrieval and visual grounding) performance from MMEB, as shown in Tab. 11. For cropping and text-only modes, we remove the visual prompt streamline. All I2T evaluations are reported at matched precision (bf16), and all results are obtained with a batch size of 1 on a single H100 GPU. While cropping and text-only modes offer better latency and memory, the performance becomes significant inferior. We further include the removal of visual prompts for I2I scenarios, which is more efficient but degrading significantly. Moreover, using uniform points is more efficient than the existing baselines.
E.7 Detailed Scores of MMEB
@@
![[Uncaptioned image]](x5.png)
![[Uncaptioned image]](x6.png)
![[Uncaptioned image]](x7.png)
![[Uncaptioned image]](x8.png)
![[Uncaptioned image]](x9.png)